please help me solve this A copy machine makes 28 copies per minute. How many copies does it make in 3 minutes and 30 seconds?

Answer by mananth(16946)

A copy machine makes 28 copies per minute. How many copies does it make in 3 minutes and 30 seconds?
A copy machine makes 28 copies per minute
For doing any calculation the units must be the same in the problem
You can either convert the time to minutes or seconds .
The question here is 3 min 30 seconds
convert to minutes
3 minutes 30 seconds = 3 1/2 minutes = 7/2 minutes
Direct variation. ie: if time increases the the number of copies made increases
1 minute 28 copies
in 3 1/2 minutes ?
7/2 * 28 = 98 copies